A car’s odometer (the meter that measures the distance traveled by the car) reads 25,943 km at the start of a trip. It the reads 26,246 km at the end of the trip. The entire trip took a total of 3.5 hours. Determine the car’s average speed in units of (a) km/h and (b) m/s.


1
Expert's answer
2021-10-19T15:12:36-0400

Given:

"x_i=25\:943\:\rm km"

"x_f=26\:246\:\rm km"

"t=3.5\:\rm hr"


The average speed

"v=\frac{x_f-x_i}{t}"

(a)

"v=\rm\frac{26\:246\: km-25\:943\: km}{3.5\: hr}=86.6\: km/hr"

(b)

"\rm 86.6\: km/hr=86.6*\frac{1000\: m}{3600\: s}=24.0\: m/s"